The demand and supply equations for a certain type of sport sandals on a selected day are found as QS = 20 + 4p and QD = 90 – 3p; The market price at the equilibrium is also found as p = $10. Thus, we can conclude that at the market equilibrium:


a. Market demand = market supply = 60 units


b. Sales revenue per day from the sale of thongs = $60


c. Slope of the demand curve = slope of the supply curve


d. All of the above are correct
Answer #1

The market equilibrium price is $10 per unit.

Demand equation is as follows -

QD = 90 - 3p

QD = 90 - (3 * 10)

QD = 90 - 30

QD = 60

The market demand at equilibrium is 60 units.

Supply equation is as follows -

QS = 20 + 4p

QS = 20 + (4 * 10)

QS = 20 + 40

QS = 60

The market supply at equilibrium is 60 units.

Thus,

At market equilibrium,

Market demand = Market supply = 60 units

Hence, the correct answer is the option (a).